OpenStudy (anonymous):

Sum the series 1x3+3x6+5x9+ ... up to n terms. Further I want to know how do we find the common difference of such a sequence (d) in order to find term k= a +(k-1)d. And the use term K to find the sum.

ganeshie8 (ganeshie8):

start by finding the general term

ganeshie8 (ganeshie8):

the \(n^{th}\) term..

OpenStudy (anonymous):

how do we find k that is the difference first the book says k is 2

ganeshie8 (ganeshie8):

what is "k" ?

OpenStudy (anonymous):

oh sorry my bad d is 2 what I get is that they did 3-1=2 or 5-3=2 to get d but how dowe know which no of the sequence to subtract for common difference if we use 6-3 it gives us 3

ganeshie8 (ganeshie8):

Got you :) I would start like this :- I notice that each term has two parts, so the \(n^{th}\) term need to look somethig like : \(part1 \times part2\)

OpenStudy (anonymous):

hmmm go on

ganeshie8 (ganeshie8):

\(part2\) is easy to figure out : its simply \(3n\), so the \(n^{th}\) term becomes : \(part1\times 3n\)

ganeshie8 (ganeshie8):

does this really look correct ? lets check it once quick

ganeshie8 (ganeshie8):

\(a_n = part1 \times 3n\) put n = 1, you wud get : \(a_1 = part1 \times 3\) put n = 2, you wud get : \(a_2 = part1 \times 6\) put n = 3, you wud get : \(a_3 = part1 \times 9\) ...

ganeshie8 (ganeshie8):

looks like we're right, it has to be \(3n\) so lets go ahead and figure out \(part1\)

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

\(part1\) has below pattern : 1, 3, 5, .... right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

does that look like a familiar sequence ?

OpenStudy (anonymous):

I have A query how do we get to know whether we have to take part 1 or part 2 to take out d?

ganeshie8 (ganeshie8):

for both parts we have a "d", but there is no "d" for the series as a whole

ganeshie8 (ganeshie8):

the "d" for first part is "3" the "d" for second part is "2"

OpenStudy (anonymous):

ssecond is 3 first is 2

ganeshie8 (ganeshie8):

good catch :) you're right !!

OpenStudy (anonymous):

;D but which one are we supposed to take?

ganeshie8 (ganeshie8):

wat do we get for \(part1\) ?

ganeshie8 (ganeshie8):

there is NO "d" for the series as whole, lets keep going :)

OpenStudy (anonymous):

Okay I think I got it if we take d is 2 then we can take first term as 1 , while if we take d as 3 we can take first term as 3.

OpenStudy (anonymous):

we form an equation and get term k (if we take d as 2) 2k-1 while if we take d as 3 Term k is: 3k

ganeshie8 (ganeshie8):

yes :) for \(part1\) : 1, 3, 5... since this is an arithmetic sequence, \(n^{th}\) term, \(a_n = 1 + (n-1)\times 2 = 1 + 2n -2 = 2n -1 \)

ganeshie8 (ganeshie8):

so, the \(n^{th}\) term of series is : \(a_n = (2n-1)(3n)\)

OpenStudy (anonymous):

we multiply both these terms to get the final term k that is 2k-1 * 3k = 6k^2 -3k

ganeshie8 (ganeshie8):

looks good

OpenStudy (anonymous):

well thank you I know the rest as to how to sum it if it is k ,k^2,k^3 etc

ganeshie8 (ganeshie8):

Sum of first \(n\) terms = \(\large \sum \limits_{k=1}^n (6k^2 - 3k)\)

ganeshie8 (ganeshie8):

yes, rest should be easy... have fun ;P

OpenStudy (anonymous):

yup